OpenStudy (anonymous):

Solve for x given ABCD AD= 10+8x BC= 16+2x

OpenStudy (anonymous):

AD and BC are equivalent as are AB and CD so 10+8x=16+2x or 10+8x-16+2x=0 Move the x's to one side of the = and the numbers to the other side for the answer.

OpenStudy (anonymous):

Slight error The equation should be 10+8x -16-2x=0

OpenStudy (anonymous):